OpenStudy (anonymous):

Use substitution to solve: int_{0}^{ln(sqrt{3})}(e^x/(1+e ^{2x}))dx

OpenStudy (anonymous):

\[\int\limits_{0}^{\ln(\sqrt3)}(e^x/(1+e^{2x}))dx\]

OpenStudy (anonymous):

let u = e^(x) so du = e^x*dx then change your limits bottom limit is e^0 = 1 top limit is ln(e^(√3)) = √3 the integrand should simplify to du/(1+x^2) which is the arctan. Then evaluate at the endpoints and subtract.

OpenStudy (anonymous):

why is the top limit sqrt(3)? shouldent it be e^(ln(sqrt3))

OpenStudy (anonymous):

well i guess that is the same thing so nvm lol ty